OpenStudy (anonymous):

Step 1: |x - 2| + 3 = 7 Step 2: |x - 2| = 7 - 3 Step 3: |x - 2| = 4 Which of the following is a correct next step to solve the equation? x + 2 = -4 -x – 2 = 4 x + 2 = 4 x - 2 = -4

OpenStudy (anonymous):

@pgpilot326 since the absolute value cant be negative would that be x+2=4?

OpenStudy (anonymous):

so think |something|=4 what can that something possibly be?

OpenStudy (anonymous):

in other words, which numbers are 4 away from 0?

OpenStudy (cwrw238):

x - 2 could be equal to 4 or -4

OpenStudy (cwrw238):

in both cases the absolute value is 4

OpenStudy (cwrw238):

so which option is correct?

OpenStudy (anonymous):

c?

OpenStudy (cwrw238):

no

OpenStudy (cwrw238):

as i said x - 2 could be equal to -4

OpenStudy (anonymous):

so x-2=-4

OpenStudy (cwrw238):

yes whats inside the modulus lines can be positive or negative

OpenStudy (cwrw238):

there are 2 roots in the solution set x = 2 and x = 6

OpenStudy (cwrw238):

sorry x = -2

OpenStudy (anonymous):

if @cwrw238 isn't explaining it to your satisfaction, let me know and I'm happy to help

OpenStudy (anonymous):

im really confused i thought the numbers inside the symbol couldnt be a negative?

OpenStudy (cwrw238):

no they can be - its the modulus lines which tell you to take the absolute value t(he positive value)

OpenStudy (anonymous):

okay thanks i kind of get it

OpenStudy (cwrw238):

| -4 | = 4 and | 4 | = 4

OpenStudy (anonymous):

so |x-2| is |x+2|

OpenStudy (the_fizicx99):

An easier way to remember this is; |x - 2| = 4 ==> x - 2 = 4 and x - 2 = -4

OpenStudy (cwrw238):

whatever 'comes out' of the modulus lines becomes positive thats not a mathematically correct way of explaining it but it might help you

OpenStudy (anonymous):

is the answer to this one x-2=-4?

OpenStudy (cwrw238):

yes

OpenStudy (cwrw238):

equally x - 2 = 4 could be the next step

OpenStudy (cwrw238):

lets do the full solution for you |x - 2| + 3 = 7 |x - 2| = 7 - 3 |x - 2| = 4 x - 2 = -4 x = -2 and x - 2 = 4 x = 6 so the solution set is {-2,6} or if you like x = -2 or x = 6

OpenStudy (anonymous):

ok that cleared it up a little bit howd the 4 become negative though?

OpenStudy (cwrw238):

well x - 2 ( the part inside the modulus lines) can be negative or positive right?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

can you help me with one more i have a idea of what the answer is

OpenStudy (cwrw238):

- I can see where you are getting a bit confused the -4 becomes negative because the x -2 can be negative if you find it easier do it this way we have either (x -2) or -(x - 2) = 4 so its x - 2 = 4 so x = 6 or -(x -2) = 4 -x + 2 = 4 -x = 2 x = -2 but the algebra is a bit harder this way

OpenStudy (anonymous):

okay thanks that makes it easier

OpenStudy (cwrw238):

right - you are making the bit inside the lines neagtive instead of the 4 - both give same answer

OpenStudy (cwrw238):

you have one more - i have about 15 minutes before i have to go

OpenStudy (anonymous):

The table shows the solution to the equation |2x – 3| - 1 = 2: Step 1 |2x – 3| = 2 + 1 Step 2 |2x – 3| = 3 Step 3 2x – 3 = 3 or 2x + 3 = 3 Step 4 2x = 6 or 2x = 0 Step 5 x = 3 or x = 0 Which is the first incorrect step

OpenStudy (anonymous):

im thinking step 3

OpenStudy (cwrw238):

yes - they have changed 2x - 3 to 2x+ 3 - that is wrong it should read 2x -3 = 3 or 2x - 3 = -3 or 2x - 3 = 3 or -(2x - 3) = 3

OpenStudy (cwrw238):

I think you've understood it now

OpenStudy (anonymous):

yes i do thanks for the help

OpenStudy (cwrw238):

yw